OpenStudy (lynn.caroline):

If f(x)=(3x+1)/5, then find: A.) f(3) and f(-7) - I got f(3)=2 and f(-7)=-4 is this right? B.the value of k if f(k)=8 C. Find a value of k such that f(k)=k

OpenStudy (shadowlegendx):

\[f(x) = (3x+1) \over 5\]

OpenStudy (loser66):

A is good now B

OpenStudy (loser66):

let the RHS =8, solve for x

OpenStudy (lynn.caroline):

I'm confused about where k came into play

OpenStudy (loser66):

just letter, ignore it.

OpenStudy (loser66):

f(x) = (3x+1)/5 f(k)=(3k+1)/5 same!!

OpenStudy (loser66):

it is not important

OpenStudy (lynn.caroline):

Okay so it makes no difference then??

OpenStudy (lynn.caroline):

So b is f(8)=5?

OpenStudy (loser66):

No, I said , let the right hand side =8, solve for x

OpenStudy (loser66):

it meant (3x+1)/5=8

OpenStudy (loser66):

now, solve for x

OpenStudy (lynn.caroline):

So x=13??

OpenStudy (pythagoras123):

Part C: \[f(k)=\frac{ 3k+1 }{ 5}=k\] \[k=\frac{ 1 }{ 2}\] (answer)

OpenStudy (lynn.caroline):

Thank you guys!!